the healthcare provider orders digoxin 250 micograma PO daily. the elixir containa 0.05 mg/mL. how much would you admininster?
Answer #1

Ordered dose=250 microgram=0.25 mg

( 1v microgram=0.001 miligram)

available dose=0.05 mg/ml

If 0.05 mg in 1 ml

then 0.25mg in X ml

therefore 0.05 * X = 0.25 * 1

X=0.25/0.05=5 ml

Therefore 5ml should be administered.